Evening shift summary (10/12)

Richard Tesarek (tesarek@fnal.gov)
Wed, 13 Oct 1999 00:07:48 -0500

Shift Summary for Evening shift, 12/Oct/99

On Shift: R.J. Tesarek, V. Prasad

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 6.5
Hours of E799 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Vacuum problems in Tevatron,
pipeline power supply
NM4AN off
Trigger problems (bananas)
Initializing



problems.

========================================================================
Runs ended this shift:

Run 14826 - BEAM_799 - 3.9M events
Run 14827 - BEAM_799 - special magnet off run
Run 14829 - BEAM_799 - 0.2M events

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: Fix pipeline (crate 12) power supply.

========================================================================
Overview of shift:
Shift started with no beam and vacuum problems in the tevatron.
Technicians "fixed" NM2SR power supply, shift crew notices current
is still low when beam returns. Beam arrives and we run smoothly
until 20:06. Install UPH009 and the KTeV gremlins awake (don't mess
with the hyperon tapes). DDD veto problems lead to pipeline resets
which fail in crate 12. BobT instantly appears 2 min after the page
and determines that a power supply is at fault. Access made to swap
the offending power supply. Access complete; we come up in NM4AN off
setting for special run. Magnet on, SDAQ problems fixed with a
call to A. Alavi-Harati. Begin run and learn of Banana problem.
J. Ladue phoned and arrives to fix the problem. Problem fixed, new
run initiated.

New XT maps loaded at end of run 14826.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14826 6 18:19 KTEVANA errors (missing FFFF word)
stream=5: RWP=9.
9 18:23 DAQDEAD alarm SEM=9.3E12.
25 18:42 DAQDEAD alarm SEM=8.9E12
20:06 Load UPH009 into tps3d2 (KTEV1)
DDD veto on, KFREND
Pipeline problems crate 8,12 (reboot)
Pipeline problems crate 12 (reboot)
20:33 Stop run access requested.
14827 21:22 Start "special" run.
21:40 discover NM4AN was off (from access?)
end special run.
14829 21:45 Initialize and start run.
22:51 Stop run due to banana problems
(last 30 spills not usable)
14830 22:56 Start run

========================================================================
Follow up Notes on Previous Shift

SDAQ: NM2SR readback is a real problem. Rick Ford looked at the history of
the readings and the values has indeed been fluctuating recently. He
called the MCR. They will check the power supply

========================================================================
Items to Follow up on Next Shift

Keep an eye on the readback of NM2SR.

Known problems
--------------
Online histos: Banana 1X-3 (sometimes 1X-2), RC10 timing,
DC planes 9&10 inefficiency in beam region,
DC4Y intrinsic efficiency low, DC3X high SOD rms,
CsI 961 comparator errors, CsI 2310 7% gain drift.
BA shows 3 channels with low gain ("weak ADCs") 2,10,73.
ETOT effic slightly low (0.998)